OpenStudy (anonymous):

18/3 = 6 is an example of division. What is 24/4? A. 6 B. 12 C. 18 D. 24

OpenStudy (anonymous):

24/4=6

OpenStudy (anonymous):

24/4=6 because 4*6=24